OpenStudy (anonymous):

write the mixed number as a improper fraction 1 7/10

OpenStudy (anonymous):



OpenStudy (anonymous):

17/10

OpenStudy (anonymous):

17 / 10

OpenStudy (anonymous):

Are you posting your entire assignment on here? -_- 1= 10/10 So 7/10 + 10/10 = 17/10